# BOWIE-ALIGN: How formation and migration histories of giant planets impact atmospheric compositions

### v2

@@ 1 Introduction
Given these complexities, it is natural to ask whether it is possible to identify a sample of exoplanets that: i) likely had different formation histories and ii) are relatively amenable to the characterisation of their atmospheric compositions. Such a sample can be used to test whether exoplanet atmosphere compositions correlate with their formation history and determine which planet formation aspects are most readily addressed through atmospheric compositions. The BOWIE+ (Bristol, Oxford, Warwick, Imperial, Exeter, +) collaboration’s BOWIE-ALIGN survey (A spectral Light Investigation into hot gas Giant origiNs, Kirk et al. (2024)) seeks to achieve this by characterising the atmospheric composition of eight hot Jupiters with JWST, four with orbits aligned with their stellar spin axes and four that are misaligned.
@@ 2 Planet formation model
To understand the evolution in composition during planet formation, we build a simple model of a planet growing inside a gas and dust disc containing various molecular species. With this model, we explore how planet composition depends on assumptions regarding the formation process and the uncertain disc environment, e.g., disc mass, temperature, dust, and planetesimal flux. The BOWIE-ALIGN programme planets orbit F-stars above the Kraft break that do not tidally align the planets. Therefore, this investigation focuses on F-star systems rather than solar analogues. Our approach closely follows existing state-of-the-art planet evolution models (Fortier et al., 2013; Booth et al., 2017; Schneider & Bitsch, 2021; Danti et al., 2023). The model is designed to be simple enough to explore a large range of parameters, investigating the variety of possible outcomes for different growth scenarios of hot gas giants. A detailed description of the model can be found in Appendix A.
The planet grows by accreting solids and gases from the surrounding disc. We adapted the description of planet growth from Bitsch et al. (2015) for planet formation via pebble accretion to include a parameterised radially dependent planetesimal accretion based on the findings in Fortier et al. (2013); a similar approach was taken by Danti et al. (2023). The planets accrete from the local environment, composed of dust and gas. For simplicity, the surface density of planetesimals is set to a constant multiple of local dust surface density. This planetesimal-to-dust mass fraction $f_{\rm pl}$ is varied between $10^{-5}-10^{0}$. Since we consider atmospheric growth of super-Earth protoplanets ($m_{\mathrm{ini}}=5~{}M_{\oplus}$) into giants, planet migration is treated using a simplified prescription dependent on the viscous speed, planet mass and disc density derived from fits to the results of the simulations by Dürmann & Kley (2015).
@@
The protoplanetary disc is modelled as a steady-state viscous disc with a Shakura-Sunyaev viscosity of $\alpha=10^{-3}$. We include dust dynamics by specifying the Stokes number, $St$, of the dust grains, which controls how fast the dust grains migrate towards the star. The dust and gas are well-coupled at low Stokes numbers ($St\ll\alpha$), moving together. We fix the $\alpha$ parameter since its effects are degenerate with the other parameters, such as disc mass and disc temperature. The composition of the gas and solids in the disc is computed by solving for the adsorption and thermal desorption of molecules from the grain surfaces, as outlined in Appendix A.1, together with the advection and diffusion of the gas and ices, as in Booth et al. (2017); Booth & Ilee (2019) or Schneider & Bitsch (2021). We do not include any chemical conversion of one species into another, instead investigating the impact of chemical evolution by exploring different abundances of the key molecules. The abundances of each of the molecular species are chosen to produce a total composition equal to the Asplund et al. (2009) solar composition.
Once the composition is specified, the model leaves us with six free parameters: the total disc mass parameterised by the mass accretion $\dot{M}_{\text{g}}$, the ratio of dust-to-gas fluxes in the disc $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$, the main size setting the drift velocities of the considered dust distribution St/$\alpha$, the fraction of mass in planetesimal relative to dust $f_{\text{pl}}$, the disc temperature, $T=T_{0}(R[au])^{-0.5}$, prescribed through the temperature at 1 au, $T_{0}$, and the initial distance of the planets to the star $r_{\text{ini}}$. All of these have been varied across the ranges listed in Tab. 1.
An example of the disc composition and resulting local C/O ratio is shown in Fig. 1 for two models with the same accretion rates of gas and solids but different Stokes numbers, $St$, and associated radial drift speeds. At the same $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$, gas phase abundances of the different models are unaffected by $St$ as it is the total mass flux, not the speed of transport, that sets the composition. The dust-to-gas ratio and ice phase abundances are lower for higher $St$, due to the steady flux of compounds between faster solid and slower gas phase. Fig. 1 also shows the familiar results that the C/O ratio of the gas and solids changes across ice lines. When radial drift is important, spikes in the ice-phase abundances appear close to ice lines because of the diffusion of gas-phase molecules back across the ice lines (Öberg & Bergin, 2016; Booth et al., 2017). Since the spikes in abundance near ice lines are a transient feature, they are not captured by our steady-state model (discussed in Sec. 4.4).
@@ 3 A zoo of modelled close-in giants
Figure 4: Probability distribution of planets for all model parameters shown in Fig.  3. The models have been given weights to produce samples with a uniform distribution on a linear scale. Thus gas dominated models are down-weighted compared to Figure 3. Orange colours show disc-migrated planets and blue colours show high-eccentricity migrated planets. The colours from light to dark incorporate [90%, 75%, 50%] of the total sample. The probability density function (PDF) of both the C/O and O/H ratio is displayed to the right and top of the figure. Only planets that purely migrated through the disc can reach C/O ratios below 0.25 and planets with C/O above 0.85 are overwhelmingly more likely to have assembled beyond 1 au if any planetesimal accretion occurs.
@@
  1. (1)
    Planets that formed with only small amounts of solids accreted through planetesimals ($f_{\text{pl}}<10^{-4}$) remain below 5×5$\times$5 × solar metallicity and above C/O0.5greater-than-or-equivalent-toabsent0.5$\gtrsim$ 0.5≳ 0.5 (marked by the dashed blue line), consistent with previous results from pebble accretion models (Booth et al., 2017; Schneider & Bitsch, 2021).
  2. (2)
    Planet that are predominately enriched by gas accretion ($\mathrm{O_{gas}}/\mathrm{O_{dust}}>1$, saturated colours) have high C/O ratios of C/O$>0.5$ throughout, due to methan and CO.
  3. (3)
@@
The difference in C/O ratio between the disc migration and high-eccentricity migration planets arises primarily due to the change in the composition of the solids across the ice lines. Particularly critical to the low C/O ratio of the disc migration planets is the accretion of solids inside $\sim 0.2$ au where the carbon grains are destroyed and, therefore, the planetesimals are then made of essentially pure silicates, which are carbon-free but contain oxygen. Since the positions of the ice lines are temperature dependent, warmer discs contribute to larger differences between the populations as the carbon destruction line (sometimes called the ‘soot line’) is further out. The different temperature models, thereby, blur the line between the populations of aligned and misaligned planets. Appendix B discusses the free parameters’ effects, including temperature, in more detail.
@@
In Fig. 4, all the models are combined into a sample density that has been weighed to factor out the logarithmic planetesimal fraction values. Here, we find two clear separation lines: only planets that purely migrated through the disc can reach C/O ratios below 0.25; planets with C/O above 0.85 are overwhelmingly more likely to have assembled beyond 1 au if significant planetesimal accretion occurs. Generally, planets assembled beyond 1 au are more metal-poor and have a higher C/O ratio, while fully migrated planets are more likely to be more metal-rich.
@@ 4.2 Carbon carriers
Figure 7: The atmosphere compositions for high and low carbon refractory abundances. Light orange and blue (x) indicate models with $75\%$ of organic carbon in refractory grains, while saturated orange and blue (+) indicate $25\%$. In the top panel, silicates enrich the atmosphere in oxygen. In the bottom panel, silicates do not contribute to the atmosphere composition.
@@ 4.4 Other considerations
In this work, we have also considered steady (time-independent) disc models. These models do not produce spikes in the gas phase abundance of molecules inside the ice lines (e.g Öberg & Bergin, 2016; Booth et al., 2017), because this is a transient effect caused by the increase in $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$ over time associated with grain growth. Given a sufficiently long time, the transient spikes seen in time-dependent models disappear as the molecules are carried inwards by gas accretion, leading to a uniformly increased abundance inside the ice line instead (see, e.g., Booth et al., 2017). Rather than using a time-dependent model, which would force us to specify when the planets form and how the dust evolves (by using the Birnstiel et al. 2012 model, for example), we simply vary $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$ to allow for an enhancement of the molecular abundances arising from radial drift. The main difference between time-dependent models and a comparable steady-state model would be a somewhat increased scatter in the C/O ratio of the planets that are not substantially enriched by solids, owing to the larger range of disc gas compositions arising from the ‘spikes’.
@@ 5 What might the BOWIE-ALIGN sample tell us about planet formation?
If the BOWIE-ALIGN survey of hot Jupiters shows differences between the disc migration (aligned) and high-eccentricity migration (misaligned) planets, as predicted by the simplified or baseline models, then this sample likely tells us something about how these planets formed or what conditions they formed in. The model distribution in Fig. 7 highlights that the current uncertainties of the physics and chemistry of disc and planet allow a wide range of formation models, that only observations can narrow down to the actual process of planet formation. Here, we summarize what the main takeaway points would be, accounting for the complications discussed in section 4. Previous observations of hot gas giants show super-solar metallicities (e.g., Alderson et al., 2023; Ahrer et al., 2023; Bean et al., 2023; Feinstein et al., 2023; Rustamkulov et al., 2023; Xue et al., 2024). This hints at a significant accretion of solids such as planetesimals, but for moderately super-solar metallicity (a few times solar), the models are degenerate because super-solar metallicity gas produced by efficient pebble accretion and drift is also possible. Together with the C/O ratios, the metallicity may lead to three scenarios with three distinct conditions.
@@ 6 Conclusions
In the BOWIE-ALIGN programme we will use JWST to compare the atmospheric compositions of four hot Jupiters that likely underwent disc migration (inferred from their aligned orbits around F stars) with four that likely migrated after the dispersal of the disc via high-eccenticity migration (inferred from their misaligned orbits). Here, we investigate how the different formation histories of the two populations can change the resulting metallicity and C/O ratio using a simple model of planet formation with a wide range of disc and accretion parameters. Many aspects of planet evolution are still uncertain, including conditions in the disc, such as the temperature, composition, disc masses, and dust properties, and the importance of different planet formation processes including planet migration and pebble accretion. Here we show how the predicted compositions of the aligned and misaligned planets vary with different disc physics assumptions.
@@ A.1 Disc model
$\mathop{}\frac{\mathrm{d}\Sigma_{\mathrm{g},i}}{\mathrm{d}t}+$ $\frac{1}{R}\frac{\partial}{\partial R}\left(R\,v_{\rm R,g}\Sigma_{\mathrm{g},i}\right)=$
@@
$\mathop{}\frac{\mathrm{d}\Sigma_{\mathrm{d},i}}{\mathrm{d}t}+$ $\frac{1}{R}\frac{\partial}{\partial R}\left(R\,v_{\rm R,d}\Sigma_{\mathrm{d},i}\right)=$
@@ A.2.3 Gas Accretion
The main part of accretion will consist of gas accretion. Westart with rocky super-Earths in which the core mass exceeds the atmosphere mass, so we consider both the hydrostatic growth phase and runaway gas accretion. We start in this regime to eliminate planets that do not reach runaway gas accretion, as their feeding zone might be too small or feeding too slow. When $m_{\mathrm{core}}<m_{\mathrm{gas}}$ the gas accretion is in the hydrostatic regime for which we use the gas accretion rates based on the Kelvin-Helmholtz time-scale from 2014Piso:

### v3

@@ 1 Introduction
Given these complexities, it is natural to ask whether it is possible to identify a sample of exoplanets that: i) likely had different formation histories and ii) are relatively amenable to the characterisation of their atmospheric compositions. Such a sample can be used to test whether exoplanet atmosphere compositions correlate with their formation history and determine which planet formation aspects are most readily addressed through atmospheric compositions. The BOWIE+ (Bristol, Oxford, Warwick, Imperial, Exeter, +) collaboration’s BOWIE-ALIGN survey (A spectral Light Investigation into hot gas Giant origiNs, (Kirk et al., 2024)) seeks to achieve this by characterising the atmospheric composition of eight hot Jupiters with JWST, four with orbits aligned with their stellar spin axes and four that are misaligned (Kirk et al., 2024).
@@ 2 Planet formation model
To understand the evolution in composition during planet formation, we build a simple model of a planet growing inside a gas and dust disc containing various molecular species. With this model, we explore how planet composition depends on assumptions regarding the formation process and the uncertain disc environment, e.g., disc mass, temperature, dust, and planetesimal flux. The BOWIE-ALIGN programme planets orbit F-stars above the Kraft break that do not tidally align the planets. Therefore, this investigation focuses on F-star systems rather than solar analogues. Our approach closely follows existing state-of-the-art planet evolution models (Fortier et al., 2013; Booth et al., 2017; Schneider & Bitsch, 2021; Danti et al., 2023). The model is designed to be simple enough to explore a large range of parameters, investigating the variety of possible outcomes for different growth scenarios of hot gas giants. In our model, we simplify the disc to a steady state within each simulation, while the range of different parameters cover different stages of the evolution of the disc at which the planets may form. A detailed description of the model can be found in Appendix A.
The planet grows by accreting solids and gases from the surrounding disc. We adapted the description of planet growth from Bitsch et al. (2015) for planet formation via pebble accretion to include a parameterised radially dependent planetesimal accretion based on the findings in Fortier et al. (2013); a similar approach was taken by Danti et al. (2023). The planets accrete from the local environment, composed of dust and gas. For simplicity, the surface density of planetesimals is set to a constant multiple of local dust surface density. This planetesimal-to-dust mass fraction $f_{\rm pl}$ is varied between $10^{-5}-10^{0}$. Especially for planet growth in the inner disc where the pebble accretion isolation mass is lower, planetesimal accretion is the dominant contributor to solid accretion to the atmosphere. Since we consider atmospheric growth of super-Earth protoplanets ($m_{\mathrm{ini}}=5~{}M_{\oplus}$) into giants, planet migration is treated using a simplified prescription dependent on the viscous speed, planet mass and disc density derived from fits to the results of the simulations by Dürmann & Kley (2015).
@@
The protoplanetary disc is modelled as a steady-state viscous disc with a Shakura-Sunyaev viscosity of $\alpha=10^{-3}$. We fix the $\alpha$ parameter since its effects are degenerate with the other parameters, such as disc mass and disc temperature. We include dust dynamics by specifying a constant Stokes number, $St$, of the dust grains, which controls how fast the dust grains migrate towards the star. At low Stokes numbers ($St\ll\alpha$), dust and gas are well-coupled, moving together. The dust-to-gas mass flux ratio is set at the outer boundary, at which all molecular species but H2 are in solid form. The steady state disc adjusts all molecular species to match the radially constant mass flux. Thus, increasing the dust-to-gas flux increases the disc’s overall metallicity. The composition of the gas and solids in the disc is computed by solving for the adsorption and thermal desorption of molecules from the grain surfaces, as outlined in Appendix A.1, together with the advection and diffusion of the gas and ices, as in Booth et al. (2017); Booth & Ilee (2019) or Schneider & Bitsch (2021). We do not include any chemical conversion of one species into another, instead investigating the impact of chemical evolution by exploring different abundances of the key molecules. The abundances of each of the molecular species are chosen to produce a total composition equal to the Asplund et al. (2009) solar composition.
Once the composition is specified, the model leaves us with six free parameters: the total disc mass parameterised by the mass accretion $\dot{M}_{\text{g}}$, the ratio of dust-to-gas fluxes in the disc $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$, the dust’s Stokes number over viscous coefficient St/$\alpha$, the fraction of mass in planetesimals relative to dust $f_{\text{pl}}$, the disc temperature, $T=T_{0}(R[\mathrm{au}])^{-0.5}$, prescribed through the temperature at 1 au, $T_{0}$, and the initial distance of the planets to the star $r_{\text{ini}}$. All of these have been varied across the ranges listed in Tab. 1.
An example of the disc composition and resulting local C/O ratio is shown in Fig. 1 for two models with the same accretion rates of gas and solids but different Stokes numbers, $St$, and associated radial drift speeds. At the same $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$, gas phase abundances of the different models are unaffected by $St$, except for very localised changes near the iceline as it is the total mass flux, not the speed of transport, that sets the composition. The dust-to-gas ratio and ice phase abundances are lower for higher $St$, due to the faster radial drift at fixed dust flux for increased Stokes numbers. Fig. 1 also shows the familiar results that the C/O ratio of the gas and solids changes across ice lines. When radial drift is important, spikes in the ice-phase abundances appear close to ice lines because of the diffusion of gas-phase molecules back across the ice lines (Öberg & Bergin, 2016; Booth et al., 2017). Since spikes in gas abundance near ice lines are a transient feature, they are not captured by our steady-state model (discussed in Sec. 4.4).
@@ 3 A zoo of modelled close-in giants
Figure 4: Probability distribution of planets for all model parameters shown in Fig.  3. The models have been given weights to produce samples with a uniform distribution of all input parameters on a linear scale. Thus gas dominated models are down-weighted compared to Figure 3. Orange colours show disc-migrated planets and blue colours show high-eccentricity migrated planets. The colours from light to dark incorporate [90%, 75%, 50%] of the total sample. The probability density function (PDF) of both the C/O and O/H ratio is displayed to the right and top of the figure. Only planets that purely migrated through the disc can reach C/O ratios below 0.25 and planets with C/O above 0.85 are overwhelmingly more likely to have assembled beyond 1 au if any planetesimal accretion occurs.
@@
  • (1)
    Planets that formed with only small amounts of solids accreted through planetesimals ($f_{\text{pl}}<10^{-4}$) remain below 5×5$\times$5 × solar metallicity and above C/O0.5greater-than-or-equivalent-toabsent0.5$\gtrsim$ 0.5≳ 0.5 (marked by the dashed grey line), consistent with previous results from pebble accretion models (Booth et al., 2017; Schneider & Bitsch, 2021).
  • (2)
    Planet that are predominately enriched by gas accretion ($\mathrm{O_{gas}}/\mathrm{O_{dust}}>1$, saturated colours) have high C/O ratios of C/O$>0.5$ throughout, due to methane and CO.
  • (3)
@@
The difference in C/O ratio between the disc migration and high-eccentricity migration planets arises primarily due to the change in the composition of the solids across the ice lines. Particularly critical to the low C/O ratio of the disc migration planets is the accretion of solids inside $\sim 0.2$ au where the carbon grains are destroyed and, therefore, the planetesimals are then made of essentially pure silicates, which are carbon-free but contain oxygen. Since the positions of the ice lines are temperature dependent, warmer discs contribute to larger differences between the populations as the carbon destruction line (sometimes called the ‘soot line’) is further out. In the regime of metal-poor planets, the small separations between icelines leads to a larger scatter in the abundance of the accreted gas which can be seen in the crosses in Fig. 3. The different temperature models, thereby, blur the line between the populations of aligned and misaligned planets. Appendix B discusses the free parameters’ effects, including temperature, in more detail.
@@
In Fig. 4, all the models are combined into a sample density that now sampled with a uniform distribution of planetesimal fraction values. Here, we find two clear separation lines: only planets that purely migrated through the disc can reach C/O ratios below 0.25; planets with C/O above 0.85 are overwhelmingly more likely to have assembled beyond 1 au if significant planetesimal accretion occurs. Generally, planets assembled beyond 1 au are more metal-poor and have a higher C/O ratio, while fully migrated planets are more likely to be more metal-rich.
@@ 4.2 Carbon carriers
Figure 7: The atmosphere compositions for high and low carbon refractory abundances. Light orange and blue (x) indicate models with $75\%$ of organic carbon in refractory grains, while /newsaturated orange and blue (o) indicate $25\%$. In the top panel, silicates enrich the atmosphere in oxygen. In the bottom panel, silicates do not contribute to the atmosphere composition.
@@ 4.4 Other considerations
In this work, we have also considered steady (time-independent) disc models. The steady condition considerably reduces the computational time and allows us to run a full range of scenarios that, through the considered different disc properties, can capture stages of the evolving disc within the complete sample. These steady disc models do not produce spikes in the gas phase abundance of molecules inside the ice lines (e.g Öberg & Bergin, 2016; Booth et al., 2017), because this is a transient effect caused by the increase in $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$ over time associated with grain growth. Given a sufficiently long time, the transient gas spikes seen in time-dependent models disappear as the molecules are carried inwards by gas accretion, leading to a uniformly increased abundance inside the ice line instead (see, e.g., Booth et al., 2017). Rather than using a time-dependent model, which would force us to specify when the planets form and how the dust evolves (by using the Birnstiel et al. 2012 model, for example), we simply vary $\dot{M}_{\text{d}}/\dot{M}_{\text{g}}$ to allow for an enhancement of the molecular abundances arising from radial drift. The main difference between time-dependent models and a comparable steady-state model would be a somewhat increased scatter in the C/O ratio of the planets that are not substantially enriched by solids, owing to the larger range of disc gas compositions arising from the ‘spikes’.
@@ 5 What might the BOWIE-ALIGN sample tell us about planet formation?
If the BOWIE-ALIGN survey of hot Jupiters shows differences between the disc migration (aligned) and high-eccentricity migration (misaligned) planets, as predicted by the simplified or baseline models, then this sample likely tells us something about how these planets formed or what conditions they formed in. The model distribution in Fig. 7 highlights that the current uncertainties of the physics and chemistry of disc and planet allow a wide range of formation models, that only observations can narrow down to the actual process of planet formation. Here, we summarize what the main takeaway points would be, accounting for the complications discussed in section 4. Previous observations of hot gas giants show super-solar metallicities (e.g., Alderson et al., 2023; Ahrer et al., 2023; Bean et al., 2023; Feinstein et al., 2023; Rustamkulov et al., 2023; Xue et al., 2024). This hints at a significant accretion of solids such as planetesimals, but for moderately super-solar metallicity (a few times solar), the models are degenerate because super-solar metallicity gas produced by efficient pebble accretion and drift is also possible. The pebbles in the models for atmospheric assembly mainly enrich the accreted gas when reaching the inner disc rather than being directly accreted as the final planet mass is several times larger than the pebble isolation mass. Together with the C/O ratios, the metallicity may lead to three scenarios with three distinct conditions.
@@ 6 Conclusions
In the BOWIE-ALIGN programme (Kirk et al., 2024) we will use JWST to compare the atmospheric compositions of four hot Jupiters that likely underwent disc migration (inferred from their aligned orbits around F stars) with four that likely migrated after the dispersal of the disc via high-eccenticity migration (inferred from their misaligned orbits). Here, we investigate how the different formation histories of the two populations can change the resulting metallicity and C/O ratio using a simple model of planet formation with a wide range of disc and accretion parameters. Many aspects of planet evolution are still uncertain, including conditions in the disc, such as the temperature, composition, disc masses, and dust properties, and the importance of different planet formation processes including planet migration and pebble accretion. Here we show how the predicted compositions of the aligned and misaligned planets vary with different disc physics assumptions.
@@ A.1 Disc model
$\frac{\partial\Sigma_{\mathrm{g},i}}{\partial t}+$ $\frac{1}{R}\frac{\partial}{\partial R}\left(R\,v_{\rm R,g}\Sigma_{\mathrm{g},i}\right)=$
@@
$\frac{\partial\Sigma_{\mathrm{d},i}}{\partial t}+$ $\frac{1}{R}\frac{\partial}{\partial R}\left(R\,v_{\rm R,d}\Sigma_{\mathrm{d},i}\right)=$
@@ A.2.3 Gas Accretion
The main part of accretion will consist of gas accretion. Westart with rocky super-Earths in which the core mass exceeds the atmosphere mass, so we consider both the hydrostatic growth phase and runaway gas accretion. We start in this regime to eliminate planets that do not reach runaway gas accretion, as their feeding zone might be too small or feeding too slow. When $m_{\mathrm{core}}<m_{\mathrm{gas}}$ the gas accretion is in the hydrostatic regime for which we use the gas accretion rates based on the Kelvin-Helmholtz time-scale from Piso & Youdin (2014):